Printer A prints 8192 character per min and printer B prints 13862 character per min. Four characters are equal to one word. Printer A starts at 7:15 am while Printer B starts at 7:29 am then at what time both will have same no of words printed.
Answer by richwmiller(17219)

It is hard to read your gibberish because you don't use capital letters nor periods consistently.
"four character are equal to one word"
rt=words
8192/4*(t+14)=13862/4*t
8192*(t+14)=13862*t yields the same result
t=20.2272 minutes = about 20 min 14 sec
7:49:14 am